# Tight Bounds for Schrödinger Potential Estimation in Unpaired Data Translation

### v2

@@ 4 Proof Insights
at $\lambda=1$, that is, ${\mathsf{KL}}(\pi^{*},\pi^{\varphi})=\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})$. In Appendix B, we prove the following general result relating variance and the cumulant generating function of a sub-exponential random variable.
Lemma 4.1.
@@
Using this lemma, we show that (see Lemma D.1)
@@
Lemma 4.2.
@@
The proof of Lemma 4.2 is postponed to Appendix C. It immediately yields that if $T$ satisfies the conditions of Theorem 1, then either the variance is sufficiently small, so that $\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)\leqslant\Upsilon(n,\delta)$, or
as desired. On the other hand, according to Lemma D.4, it holds that
Appendix A Proof of Theorem 1
Step 1: uniform concentration.  To start with, we would like to note that Assumptions 3 and 4 imply that (see Lemma E.1)
On the other hand, according to Lemma E.2, under Assumptions 2 and 4, we have
for all $\varphi\in\Phi$. This means that we can exploit properties of sub-exponential random variables (including concentration bounds) during analysis of the empirical risk minimizer $\widehat{\varphi}$. In particular, Lemma F.1 states that, under the assumptions of Theorem 1, there exists
$T_{0}\lesssim\frac{1}{b}\log\left(d+LR^{2}+\frac{Ld^{2}}{b}+M\right)$
such that for any $\delta\in(0,1)$ and any $T\geqslant T_{0}$ we have
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$ $\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
and
$\left|\frac{1}{n}\sum\limits_{j=1}^{n}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}-\mathbb{E}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\right|$ $\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
with probability at least $(1-\delta)$ simultaneously for all $\varphi\in\Phi$. We provide the proof of Lemma F.1 in Appendix F.1. It relies on Bernstein’s inequality for random variables with finite $\psi_{1}$-norms (see, for example, [Lecué and Mitchell, 2012, Proposition 5.2]) and the $\varepsilon$-net argument. This approach is quite standard. A similar strategy was used by Puchkin et al. [2025].
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\big(\varphi(Y_{i})-\varphi(Y_{i})\big)-\mathbb{E}\big(\varphi(Y_{1})-\varphi(Y_{1})\big)\right|\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi(Y_{1})-\varphi(Y_{1})\big)\Upsilon(n,\delta)}{(1\vee L/b)(M+d)}}+\Upsilon(n,\delta)$ (17)
and
$\left|\frac{1}{n}\sum\limits_{j=1}^{n}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}-\mathbb{E}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right|\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\Upsilon(n,\delta)}{(1\vee L/b)(M+d)}}+\Upsilon(n,\delta)$ (18)
simultaneously for all $\varphi\in\Phi$.
can be controlled by a sub-quadratic function of ${\mathsf{KL}}(\pi^{*},\pi^{\varphi})$. This is the key ingredient in derivation of rates of convergence possibly faster than ${\mathcal{O}}(n^{-1/2})$. We would like to note that in Puchkin et al. [2025], the authors used properties of log-densities with subquadratic growth for a similar purpose (see their Lemmata C.2 and C.3). Unfortunately, we cannot apply them in our setup, because we do not have access to samples from $\pi^{*}$, we are given only samples drawn from the marginals of $\pi^{*}$. This makes us to use another strategy. Let us observe that for any $\varphi\in\Phi$
at $\lambda=1$, that is, ${\mathsf{KL}}(\pi^{*},\pi^{\varphi})=\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})$. In Appendix B, we prove the following general result relating variance and the cumulant generating function of a sub-exponential random variable.
Lemma A.1 (restatement of Lemma 4.1).
it holds that
Based on this result, in Appendix D.1 we prove Lemma D.1 claiming that
$\mathbb{E}_{X_{0}\sim\rho_{0}}\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,|\,X_{0}\big)\lesssim\left(1\vee\frac{L}{b}\right)(M+d)\;{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)\left(1\vee\log\frac{(1\vee L/b)(M+d)}{{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)}\right).$
However, this bound is not enough, because $\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)$ is always not smaller than the conditional variance in general. Fortunately, we can use properties of the Ornstein-Uhlenbeck process (in particular, Lemma B.3 from Puchkin et al. [2025]) to derive the following result.
Lemma A.2 (restatement of Lemma 4.2).
Under the assumptions of Theorem 1, for any $\omega>0$ there exists
The proof of Lemma 4.2 is postponed to Appendix C. Taking $\omega=\Upsilon(n,1)$, we immediately obtain that there is
@@ Appendix A Proof of Theorem 1
$\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)\lesssim\Upsilon(n,\delta)+\left(1\vee\frac{L}{b}\right)(M+d)\;{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)\left(1\vee\log\frac{(1\vee L/b)(M+d)}{{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)}\right).$ (20)
On the other hand, according to Lemma D.4, it holds that
$\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)$ $\lesssim\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$
@@
This reveals a dichotomy. If $T$ is sufficiently large, then either the right-hand sides of (17) and (18) are of order $\Upsilon(n,\delta)$ or we have that the loss functions of interest meet Bernstein-type conditions (20) and (A) allowing us to derive faster rates of convergence.
Step 3: generalization error bound.  Bringing together (17), (18), (20), and (A), we obtain that, for any $\delta\in(0,1)$, with probability at least $(1-\delta)$ it holds that
@@
${\mathsf{KL}}(\pi^{*},\widehat{\pi})-\Delta\lesssim\sqrt{\Upsilon(n,\delta)\Delta\left(1\vee\log\frac{(1\vee L/b)(M+d)}{\Delta}\right)}+\Upsilon(n,\delta)\left(1\vee\log\frac{(1\vee L/b)(M+d)}{\Delta}\right)$
@@
stand for the corresponding Orlicz norm. Then, according to Lemma F.5, the derivatives of the cumulant generating function $g(\lambda)$ can be expressed as follows:
@@
It is straightforward to observe that $g^{\prime}(0)=0$ and that $g^{\prime\prime}(0)=\mathrm{Var}(\xi)\leqslant 4\|\xi-\mathbb{E}\xi\|_{\psi_{1}}^{2}$ due to Lemma F.8. Let us elaborate on $g^{\prime\prime\prime}(\lambda)$. For this purpose, let us introduce
@@
In view of Lemma F.8 and Lemma F.6, we have
@@
$\left(\mathsf{P}_{\lambda}\left|\xi-\mathbb{E}\xi\right|^{2+1/\varepsilon}\right)^{\varepsilon/(1+2\varepsilon)}\leqslant 2^{\varepsilon/(1+2\varepsilon)}\left(3+\frac{1}{\varepsilon}\right)\|\xi-\mathbb{E}\xi\|_{\psi_{1}(\mathsf{P}_{\lambda})}\leqslant 2^{\varepsilon/(1+2\varepsilon)}\left(3+\frac{1}{\varepsilon}\right)\frac{\|\xi-\mathbb{E}\xi\|_{\psi_{1}}}{1-\lambda\|\xi-\mathbb{E}\xi\|_{\psi_{1}}}.$
@@
Combining this bound with (B), we obtain that
@@
Appendix C Proof of Lemma 4.2
Let us fix an arbitrary $\omega>0$ and consider a log-potential $\varphi\in\Phi$ satisfying the inequality $\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)\geqslant\omega$. Also, let $T\gtrsim\log(d)/b$ be such that ${\mathcal{K}}(T)\leqslant\sqrt{2}$, where the function ${\mathcal{K}}(t)$ is defined in (31). The proof of Lemma 4.2 relies on Lemma D.2, which provides an upper bound on the variance of $\mathbb{E}\left[\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}\right]$. Due to the mixing of the diffusion process $X^{*}$, it is not surprising that $\mathrm{Var}\left(\mathbb{E}\left[\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}\right]\right)$ tends to zero as $T$ grows. To be more precise, Lemma D.2 claims that
@@
$\mathrm{Var}\left(\mathbb{E}\left[\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}\right]\right)$ $\lesssim(M^{2}\vee 1)de^{-2bT}\left(\frac{1}{\omega}\vee\frac{1}{\omega^{2}}\right)^{1-1/{\mathcal{K}}^{2}(T)}\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$
@@
$(M^{2}\vee 1)\,de^{-2bT}\exp\left\{{\mathcal{O}}\left(\sqrt{d}\,e^{-bT}\log\left(1\vee\frac{1}{\omega}\right)+\sqrt{d}\left(M\vee 1\vee e^{-bT}\log\left(\widetilde{M}\vee LR^{2}\vee(Ld^{2}/b)\right)\right)\right)\right\}\leqslant 1/2.$
@@
Appendix D Log-potential variance and conditional variance bounds
In this section, we elaborate on variance bounds. The presented results is a key argument in derivation of the rates of convergence, which are possibly faster than ${\mathcal{O}}(n^{-1/2})$. We use Lemma D.2 in the proof of Lemma 4.2 and then combine the results of Lemma 4.2 and Lemma D.1 to derive an important intermediate result in the proof of Theorem 1, the inequality (20).
Lemma D.1.
@@ Appendix D Log-potential variance and conditional variance bounds
$\lesssim{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)+\left(1\vee\frac{L}{b}\right)\left(d+M+R^{2}e^{-2bT}\right){\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)\log\left(1+\frac{(1\vee L/b)(d+M+R^{2}e^{-2bT})}{{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)}\right).$
@@
The proofs of Lemma D.1 and Lemma D.2 are moved to Appendices D.1 and D.2, respectively. While Lemma D.1 is an almost immediate consequence of Lemma 4.1 and the representation (19), the proof of Lemma D.2 relies on the following property of the Ornstein-Uhlenbeck operator.
Lemma D.3.
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant F_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
We provide the proof of Lemma D.3 in Appendix D.3. Finally, in Appendix D.4 we prove the following counterpart of Lemma D.1 for $\log\big({\mathcal{T}}_{T}[e^{\varphi}](X_{0})/{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})\big)$.
Lemma D.4.
@@
Similarly to Lemmata D.1 and D.2, we apply Lemma D.4 on the second step of the proof of Theorem 1 to derive the inequality (A).
D.1 Proof of Lemma D.1
@@
$G(\lambda,x_{0})=\log\mathbb{E}\left[\exp\left\{\lambda\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)-\lambda\mathbb{E}\left[\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}=x_{0}\right]\right\}\,\Big|\,X_{0}=x_{0}\right],$
@@
Then, according to Lemma 4.1,
@@
$=\frac{52e^{5}\Psi^{2}(x_{0})}{\mathrm{Var}\big(\varphi(X_{T})-\varphi^{*}(X_{T})\,|\,X_{0}=x_{0}\big)}\left(1\vee 4e\Psi(x_{0})\log\frac{52e^{5}\Psi^{2}(x_{0})}{\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,|\,X_{0}=x_{0}\big)}\right).$
Then Lemma F.9 implies that
@@
Applying Lemma E.3, we obtain that
@@
$\lesssim\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})+\left(1\vee\frac{L}{b}\right)\left(d+M+R^{2}e^{-2bT}\right)\mathbb{E}_{X_{0}\sim\rho_{0}}\left[G(1,X_{0})\log\left(1+\frac{(1\vee L/b)(d+M+R^{2}e^{-2bT})}{G(1,X_{0})}\right)\right].$
@@
$\lesssim\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})+\left(1\vee\frac{L}{b}\right)\left(d+M+R^{2}e^{-2bT}\right)\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})\log\left(1+\frac{(1\vee L/b)(d+M+R^{2}e^{-2bT})}{\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})}\right)$
$={\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)+\left(1\vee\frac{L}{b}\right)\left(d+M+R^{2}e^{-2bT}\right){\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)\log\left(1+\frac{(1\vee L/b)(d+M+R^{2}e^{-2bT})}{{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)}\right).$
@@
$M\leqslant-\mathbb{E}\varphi(X_{T}^{*})\leqslant L\mathbb{E}\left\|\Sigma^{-1/2}(X_{T}^{*}-m)\right\|^{2}+M=L\mathrm{Tr}\big(\Sigma^{-1}\mathrm{Var}(X_{T}^{*})\big)+L\left\|\Sigma^{-1/2}(\mathbb{E}X_{T}^{*}-m)\right\|^{2}+M$

### v3

@@ B.1 Proof Insights
at $\lambda=1$, that is, ${\mathsf{KL}}(\pi^{*},\pi^{\varphi})=\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})$. In Appendix B.3, we prove the following general result relating variance and the cumulant generating function of a sub-exponential random variable.
Lemma B.1.
@@
Using this lemma, we show that (see Lemma C.1)
@@
Lemma B.2.
Under the assumptions of Theorem 1, for any $\omega>0$ there exists
$\mathrm{Var}\left(\mathbb{E}\left[\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}\right]\right)$ $\leqslant\frac{1}{2}\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$
$\leqslant\mathbb{E}\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,|\,X_{0}\big).$
The proof of Lemma B.2 is postponed to Appendix B.4. It immediately yields that if $T$ satisfies the conditions of Theorem 1, then either the variance is sufficiently small, so that $\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)\leqslant\Upsilon(n,\delta)$, or
as desired. On the other hand, according to Lemma C.4, it holds that
B.2 Rigorous proof
Step 1: uniform concentration.  To start with, we would like to note that Assumptions 3 and 4 imply that (see Lemma D.1)
On the other hand, according to Lemma D.2, under Assumptions 2 and 4, we have
for all $\varphi\in\Phi$. This means that we can exploit properties of sub-exponential random variables (including concentration bounds) during analysis of the empirical risk minimizer $\widehat{\varphi}$. In particular, Lemma E.1 states that, under the assumptions of Theorem 1, there exists
$T_{0}\lesssim\frac{1}{b}\log\left(d+LR^{2}+\frac{Ld^{2}}{b}+M\right)$
such that for any $\delta\in(0,1)$ and any $T\geqslant T_{0}$ we have
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$
$\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
and
$\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
with probability at least $(1-\delta)$ simultaneously for all $\varphi\in\Phi$. We provide the proof of Lemma E.1 in Appendix E.1. It relies on Bernstein’s inequality for random variables with finite $\psi_{1}$-norms (see, for example, (Lecué & Mitchell, 2012, Proposition 5.2)) and the $\varepsilon$-net argument. This approach is quite standard. A similar strategy was used by Puchkin et al. (2025).
$\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi(Y_{1})-\varphi(Y_{1})\big)\Upsilon(n,\delta)}{(1\vee L/b)(M+d)}}+\Upsilon(n,\delta)$ (25)
and
$\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\Upsilon(n,\delta)}{(1\vee L/b)(M+d)}}+\Upsilon(n,\delta)$ (26)
simultaneously for all $\varphi\in\Phi$.
can be controlled by a sub-quadratic function of ${\mathsf{KL}}(\pi^{*},\pi^{\varphi})$. This is the key ingredient in derivation of rates of convergence possibly faster than ${\mathcal{O}}(n^{-1/2})$. We would like to note that in Puchkin et al. (2025), the authors used properties of log-densities with subquadratic growth for a similar purpose (see their Lemmata C.2 and C.3). Unfortunately, we cannot apply them in our setup, because we do not have access to samples from $\pi^{*}$, we are given only samples drawn from the marginals of $\pi^{*}$. This makes us to use another strategy. Let us observe that for any $\varphi\in\Phi$
at $\lambda=1$, that is, ${\mathsf{KL}}(\pi^{*},\pi^{\varphi})=\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})$. In Appendix B.3, we prove the following general result relating variance and the cumulant generating function of a sub-exponential random variable.
Lemma B.3 (restatement of Lemma B.1).
it holds that
Based on this result, in Appendix C.1 we prove Lemma C.1 claiming that
However, this bound is not enough, because $\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)$ is always not smaller than the conditional variance in general. Fortunately, we can use properties of the Ornstein-Uhlenbeck process (in particular, Lemma B.3 from Puchkin et al. (2025)) to derive the following result.
Lemma B.4 (restatement of Lemma B.2).
@@ B.2 Rigorous proof
The proof of Lemma B.2 is postponed to Appendix B.4. Taking $\omega=\Upsilon(n,1)$, we immediately obtain that there is
@@
$\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$ $\lesssim\Upsilon(n,\delta)+\left(1\vee\frac{L}{b}\right)(M+d)$
$\quad\cdot{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)\left(1\vee\log\frac{(1\vee L/b)(M+d)}{{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)}\right).$ (28)
On the other hand, according to Lemma C.4, it holds that
$\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)$
$\lesssim\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$
@@
This reveals a dichotomy. If $T$ is sufficiently large, then either the right-hand sides of (B.2) and (B.2) are of order $\Upsilon(n,\delta)$ or we have that the loss functions of interest meet Bernstein-type conditions (B.2) and (B.2) allowing us to derive faster rates of convergence.
Step 3: generalization error bound.  Bringing together (B.2), (B.2), (B.2), and (B.2), we obtain that, for any $\delta\in(0,1)$, with probability at least $(1-\delta)$ it holds that
@@
$+\Upsilon(n,\delta)\left(1\vee\log\frac{(1\vee L/b)(M+d)}{\Delta}\right)$
@@
stand for the corresponding Orlicz norm. Then, according to Lemma E.5, the derivatives of the cumulant generating function $g(\lambda)$ can be expressed as follows:
@@
It is straightforward to observe that $g^{\prime}(0)=0$ and that $g^{\prime\prime}(0)=\mathrm{Var}(\xi)\leqslant 4\|\xi-\mathbb{E}\xi\|_{\psi_{1}}^{2}$ due to Lemma E.8. Let us elaborate on $g^{\prime\prime\prime}(\lambda)$. For this purpose, let us introduce
@@
In view of Lemma E.8 and Lemma E.6, we have
@@
$\left(\mathsf{P}_{\lambda}\left|\xi-\mathbb{E}\xi\right|^{2+1/\varepsilon}\right)^{\varepsilon/(1+2\varepsilon)}$ $\leqslant 2^{\varepsilon/(1+2\varepsilon)}\left(3+\frac{1}{\varepsilon}\right)\|\xi-\mathbb{E}\xi\|_{\psi_{1}(\mathsf{P}_{\lambda})}$
$\leqslant 2^{\varepsilon/(1+2\varepsilon)}\left(3+\frac{1}{\varepsilon}\right)\frac{\|\xi-\mathbb{E}\xi\|_{\psi_{1}}}{1-\lambda\|\xi-\mathbb{E}\xi\|_{\psi_{1}}}.$
@@
Combining this bound with (B.3), we obtain that
@@
B.4 Proof of Lemma B.2
Let us fix an arbitrary $\omega>0$ and consider a log-potential $\varphi\in\Phi$ satisfying the inequality $\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)\geqslant\omega$. Also, let $T\gtrsim\log(d)/b$ be such that ${\mathcal{K}}(T)\leqslant\sqrt{2}$, where the function ${\mathcal{K}}(t)$ is defined in (39). The proof of Lemma B.2 relies on Lemma C.2, which provides an upper bound on the variance of $\mathbb{E}\left[\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}\right]$. Due to the mixing of the diffusion process $X^{*}$, it is not surprising that $\mathrm{Var}\left(\mathbb{E}\left[\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}\right]\right)$ tends to zero as $T$ grows. To be more precise, Lemma C.2 claims that
@@
$\lesssim(M^{2}\vee 1)de^{-2bT}\left(\frac{1}{\omega}\vee\frac{1}{\omega^{2}}\right)^{1-1/{\mathcal{K}}^{2}(T)}\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)$
@@
$(M^{2}\vee 1)\,de^{-2bT}\exp\left\{{\mathcal{O}}\left(\sqrt{d}\,e^{-bT}\log\left(1\vee\frac{1}{\omega}\right)\right)\right\}$
$\quad\cdot\exp\left\{{\mathcal{O}}\left(\sqrt{d}\left(M\vee 1\vee e^{-bT}\log\left(\widetilde{M}\vee LR^{2}\vee(Ld^{2}/b)\right)\right)\right)\right\}\leqslant 1/2.$
@@
Appendix C Log-potential variance and conditional variance bounds
In this section, we elaborate on variance bounds. The presented results is a key argument in derivation of the rates of convergence, which are possibly faster than ${\mathcal{O}}(n^{-1/2})$. We use Lemma C.2 in the proof of Lemma B.2 and then combine the results of Lemma B.2 and Lemma C.1 to derive an important intermediate result in the proof of Theorem 1, the inequality (B.2).
Lemma C.1.
@@ Appendix C Log-potential variance and conditional variance bounds
$\mathbb{E}_{X_{0}\sim\rho_{0}}\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}\big)$ $\lesssim{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)$
$\quad+\left(1\vee\frac{L}{b}\right)\left(d+M+R^{2}e^{-2bT}\right){\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)$
$\quad\cdot\log\left(1+\frac{(1\vee L/b)(d+M+R^{2}e^{-2bT})}{{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)}\right).$
@@
The proofs of Lemma C.1 and Lemma C.2 are moved to Appendices C.1 and C.2, respectively. While Lemma C.1 is an almost immediate consequence of Lemma B.1 and the representation (27), the proof of Lemma C.2 relies on the following property of the Ornstein-Uhlenbeck operator.
Lemma C.3.
@@
$\mathrm{Var}\left(\frac{{\mathcal{T}}_{T}[fe^{\varphi^{*}}](X_{0})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})}\right)$ $\leqslant F_{\max}^{2}\left(\frac{{\mathcal{T}}_{\infty}[fe^{\varphi^{*}}]}{F_{\max}{\mathcal{T}}_{\infty}[e^{\varphi^{*}}]}\right)^{2/{\mathcal{K}}(T)}$
$\quad\cdot\exp\left\{2M\big({\mathcal{K}}(T)-1\big)+2\overline{{\mathcal{A}}}(T)\left({\mathcal{K}}(T)+\frac{1}{{\mathcal{K}}(T)}\right)\right\}$
@@
We provide the proof of Lemma C.3 in Appendix C.3. Finally, in Appendix C.4 we prove the following counterpart of Lemma C.1 for $\log\big({\mathcal{T}}_{T}[e^{\varphi}](X_{0})/{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})\big)$.
Lemma C.4.
@@
Similarly to Lemmata C.1 and C.2, we apply Lemma C.4 on the second step of the proof of Theorem 1 to derive the inequality (B.2).
C.1 Proof of Lemma C.1
@@
$G(\lambda,x_{0})$ $=\log\mathbb{E}\left[\exp\left\{\lambda\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\big)\right\}\,\Big|\,X_{0}=x_{0}\right]$
$\quad-\lambda\mathbb{E}\left[\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,\big|\,X_{0}=x_{0}\right],$
@@
Then, according to Lemma B.1,
@@
$=\frac{52e^{5}\Psi^{2}(x_{0})}{\mathrm{Var}\big(\varphi(X_{T})-\varphi^{*}(X_{T})\,|\,X_{0}=x_{0}\big)}$
$\quad\cdot\left(1\vee 4e\Psi(x_{0})\log\frac{52e^{5}\Psi^{2}(x_{0})}{\mathrm{Var}\big(\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\,|\,X_{0}=x_{0}\big)}\right).$
Then Lemma E.9 implies that
@@
Applying Lemma D.3, we obtain that
@@
$\lesssim\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})+\left(1\vee\frac{L}{b}\right)\left(d+M+R^{2}e^{-2bT}\right)$
$\quad\cdot\mathbb{E}_{X_{0}\sim\rho_{0}}\left[G(1,X_{0})\log\left(1+\frac{(1\vee L/b)(d+M+R^{2}e^{-2bT})}{G(1,X_{0})}\right)\right].$
@@
$\lesssim\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})+\left(1\vee\frac{L}{b}\right)\left(d+M+R^{2}e^{-2bT}\right)\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})$
$\quad\cdot\log\left(1+\frac{(1\vee L/b)(d+M+R^{2}e^{-2bT})}{\mathbb{E}_{X_{0}\sim\rho_{0}}G(1,X_{0})}\right)$
$={\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)+\left(1\vee\frac{L}{b}\right)\left(d+M+R^{2}e^{-2bT}\right){\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)$
$\quad\cdot\log\left(1+\frac{(1\vee L/b)(d+M+R^{2}e^{-2bT})}{{\mathsf{KL}}\left(\pi^{*},\pi^{\varphi}\right)}\right).$
@@
$-M\leqslant-\mathbb{E}\varphi(X_{T}^{*})$ $\leqslant L\mathbb{E}\left\|\Sigma^{-1/2}(X_{T}^{*}-m)\right\|^{2}+M$
$=L\mathrm{Tr}\big(\Sigma^{-1}\mathrm{Var}(X_{T}^{*})\big)+L\left\|\Sigma^{-1/2}(\mathbb{E}X_{T}^{*}-m)\right\|^{2}+M$